OpenStudy (anonymous):

What is the range of the function y = -x2 + 1? i really need help please

OpenStudy (anonymous):

the very smallest \(x^2\) can be is zero because if you square a non zero number it is always positive that means the very LARGEST \(-x^2\) can be is zero

OpenStudy (anonymous):

so now what is the very largest \(-x^2+1\) can be?

OpenStudy (anonymous):

if it is not obvious, let me know

OpenStudy (anonymous):

o.o......uhh, 1? I really have no idea. im 100% dislexic in algebra and my only two teachers been the worst in the entire school.

OpenStudy (anonymous):

yeah it is one, because... well because if the largest \(-x^2\) can be is \(0\) then the largest \(-x^2+1\) can be is \(0+1=1\)

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

now that should answer the question, because the range is all possible \(y\) values

OpenStudy (anonymous):

\(y\) cannot be any larger than 1, so the range can be written either as \(y\leq 1\) or \[(-\infty,1]\] depending on how you are supposed to write your answer

OpenStudy (anonymous):

Im sorry for wating your time. But thank you.

OpenStudy (anonymous):

btw i knew right away you had a lousy math teacher because of the way the question was asked \[y=-x^2+1\] is not a FUNCTION it in and EQUATION

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

not a waste at all, glad to help and i hope it was more or less clear

OpenStudy (anonymous):

you can make it a function by writing \[\{(x,y):y=-x^2+1\}\] or more commonly by saying "let \(f(x)=-x^2+1\) "